OpenStudy (anonymous):

factorization of polynomials: 16u^8 - 8u^4 + 1

OpenStudy (anonymous):

since the degree is of the 1st term is just the double of the 1nd term... we can reduce it to quadratic form...

OpenStudy (anonymous):

then apply quadratic formula...

OpenStudy (anonymous):

16u^8 - 8u^4 + 1 =(4u^4)^2 -2*(4u^4)(1) +1 =(4u^4 - 1)^2

OpenStudy (anonymous):

=(4u^4 - 1)^2 = ( (2u^2)^2 - 1) ^2 = ( (2u^2)- 1) * (2u^2) + 1) )^2

OpenStudy (anonymous):

what happens to the plus one? @matricked